### v1

# InterFormer: Towards Effective Heterogeneous Interaction Learning for Click-Through Rate Prediction

Abstract.
A clear and well-documented  document is presented as an article formatted for publication by ACM in a conference proceedings or journal publication. Based on the “acmart” document class, this article presents and explains many of the common variations, as well as many of the formatting elements an author may use in the preparation of the documentation of their work.
Abstract.
Click-through rate (CTR) prediction, which predicts the probability of a user clicking an ad, is a fundamental task in recommender systems. The emergence of heterogeneous information, such as user profile and behavior sequences, depicts user interests from different aspects. A mutually beneficial integration of heterogeneous information is the cornerstone towards the success of CTR prediction. However, most of the existing methods suffer from two fundamental limitations, including (1) insufficient inter-mode interaction due to the unidirectional information flow between modes, and (2) aggressive information aggregation caused by early summarization, resulting in excessive information loss. To address the above limitations, we propose a novel module named InterFormer to learn heterogeneous information interaction in an interleaving style. To achieve better interaction learning, InterFormer enables bidirectional information flow for mutually beneficial learning across different modes. To avoid aggressive information aggregation, we retain complete information in each data mode and use a separate bridging arch for effective information selection and summarization. Our proposed InterFormer achieves state-of-the-art performance on three public datasets and a large-scale industrial dataset.
Recommendation, CTR Prediction, Heterogeneous Information
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Click-through rate (CTR) prediction, which aims to predict the probability of a user clicking an ad or item, is the fundamental task for various applications such as online advertising and recommender systems (Song et al., 2019; Lyu et al., 2020; Zhang et al., 2022, 2024). The quality of CTR prediction significantly influences the company revenue and user experience, drawing extensive attention from both academia and industry (Zhou et al., 2018, 2019; Lyu et al., 2020). For example, in online ad bidding, accurate CTR prediction helps advertisers optimize their bids and target the most receptive audiences. In content recommendation, accurate CTR prediction enables platforms to suggest more relevant content to users.
To achieve better CTR prediction, it is crucial to capture the user interests in the evolving environment (Zhou et al., 2018; Lyu et al., 2020; Wang et al., 2019). The abundance of heterogeneous information presents both opportunities and challenges. On the one hand, heterogeneous information depicts user interests from different aspects, providing diverse context (Zhang et al., 2017). For instance, global information, e.g., user profile and context features, offers a static view on general user interests, while user behavior sequences provide substantial information for modeling dynamic user interests (Wang et al., 2019). On the other hand, the heterogeneous nature of the data requires different modeling approaches and careful integration across different information modes (Zhang et al., 2017). For example, while modeling interactions among global information is critical to personalized recommendation (Rendle, 2010; Lian et al., 2018; Wang et al., 2021b), capturing sequential dependencies is the major focus for user behavior modeling (Sun et al., 2019; Chen et al., 2019).
Most of the existing CTR prediction models fall into two categories, including non-sequential models and sequential models. Non-sequential models focus on learning informative embeddings through feature interaction via inner-product (Lian et al., 2018; Sun et al., 2021), MLP (Wang et al., 2017, 2021b) and deep structured semantic model (Huang et al., 2013; Elkahky et al., 2015), but ignore the sequential information in user behaviors. Sequential models, in the contrast, employ additional modules, e.g., CNN (Tang and Wang, 2018), RNN (Sun et al., 2019; Zhou et al., 2018) and Attention modules (Lyu et al., 2020; Zhou et al., 2019; Zhai et al., 2024), to capture the sequential dependencies in user behaviors. Promising as it might be, existing sequential methods mostly employ a unidirectional information flow, where global information is used to guide sequence learning, while the reverse information flow from sequence to global information is largely ignored, hence suffering from insufficient inter-mode interaction. For example, global information often captures long-term interests, while sequence information reveals momentary interests, such as a sudden focus on a specific category of products, which can enhance the global context with immediate preference. Besides, due to the computational challenges of performing interaction learning among numerous global features and lengthy sequences, aggressive feature aggregation, e.g., sequence summation (Zhou et al., 2018), pooling (Xiao et al., 2020), and concatenation (Zhou et al., 2019), is often performed at early stages, inevitably leading to excessive information loss.
In light of the above limitations, we propose a novel heterogeneous interaction learning module named InterFormer, whose ideas are two-fold. To avoid insufficient inter-mode interaction, we enable bidirectional information flows between different modes, such that global and sequence learning are performed in an interleaving style. Specifically, to learn context-aware sequence embeddings, global sumamrization guides sequence learning via Personalized FFN (pFFN) and Multihead Attention (MHA) (Vaswani et al., 2017). To learn behavior-aware global embeddings, sequence summarization instructs global learning via an interaction module. To mitigate aggressive information aggregation, we adopt MHA for effective information selection, based on which the one-to-one mappings between input and output tokens are retained till the final interaction. Note that our framework is compatible with various interaction learning models like DCNv2 (Wang et al., 2021b), DHEN (Zhang et al., 2022), etc.
The main contributions of this paper are summarized as follows:
  • Challenges. We identify two key bottlenecks of heterogeneous interaction learning, namely insufficient inter-mode interaction and aggressive information aggregation.
  • Model Design. We propose a novel heterogeneous interaction learning framework named InterFormer for effective feature interaction and selective information aggregation. To our best knowledge, the proposed InterFormer is the first model to address the mutual benefits in heterogeneous interaction learning.
  • Experiments and Analysis. We carry out extensive experiments on the proposed InterFormer with up to 0.14% AUC improvement on benchmark datasets and 0.15% Normalized Entropy (NE) gain on internal large-scale dataset. Besides, InterFormer exhibits promising scaling results in both feature scaling and model scaling.
The rest of the paper is organized as follows. Section 2 briefly reviews the recent works on interaction learning. Section 3 summarizes the preliminaries, and section 4 introduces our proposed InterFormer. Extensive experiments and analyses are carried out in Section 5. We conclude our paper in Section 6.
2. Related Works
In this section, we briefly review the related works on feature interaction, including non-sequential and sequential methods.
2.1. Non-Sequential Methods
The vast majority of non-sequential models is built upon the idea of Factorization Machine (FM) (Rendle, 2010; Lian et al., 2018; Sun et al., 2021), which models the user-item interaction by linearly combining their low-dimensional embeddings (Zhang et al., 2019). (Rendle, 2010) is the very first FM model to capture pairwise interactions. To model high-order interactions, different methods combine FMs with deep neural networks, where FMs learn low-order interaction via pairwise operation and neural networks learn high-order interactions via deep architectures. For example, MLP (Lian et al., 2018; Yang et al., 2017; Wang et al., 2021b; Sun et al., 2021; Zhou et al., 2020) captures high-order interactions via dense connections between features, and Attention mechanism (Song et al., 2019; Xiao et al., 2017; Xin et al., 2019) learns more complex embeddings through linear combination. These deep learning-based approaches enable end-to-end training without hand-craft feature designs, and are capable of handling heterogeneous information like text, image and video (Zhang et al., 2019). Besides, recent works address the scaling law in recommendation, where DHEN (Zhang et al., 2022) ensembles multiple interaction modules and Wukong (Zhang et al., 2024) stacks FMs to learn a hierarchy of interactions. Promising as they are, non-sequential models fail to capture the sequential dependencies in user behaviors, resulting in suboptimal solutions.
2.2. Sequential Methods
With the recent emergence of sequence information, e.g., user interaction history, in recommender systems, extensive efforts have been made to capture the evolving user interest. A key challenge behind sequential methods is to combine the sequential information with the heterogeneous global information in a mutually beneficial manner. Markov models (Shani et al., 2005; He and McAuley, 2016a; Yang et al., 2020) consider sequential data as a stochastic process over discrete random variables, but the oversimplified Markovian property limits the model capability in capturing long-term sequential dependencies (Quadrana et al., 2018). To model the long-term dependencies, RNN and Attention mechanism are employed as the backbone module for many sequential methods. For example, to extract the evolving user interest, various attention-based networks are designed (Zhou et al., 2019; Lyu et al., 2020), and Transformer architecture (Devlin et al., 2018; Vaswani et al., 2017) is adopted for sequential modeling (Sun et al., 2019; Chen et al., 2019). Besides, to model multi-faceted user interests, (Xiao et al., 2020; Han et al., 2024) propose to capture multiple user interests from multi-behavior sequences. More recently, TransAct (Xia et al., 2023) adopt a hybrid ranking model to combine real-time user actions for immediate preference and batch user representations for long-term interests. LiRank (Borisyuk et al., 2024) improves CTR prediction at LinkedIn by ensembling multiple interaction modules, accelerated by quantization and vocabulary compression. CARL (Chen et al., 2024) achieves fast inference on large-scale recommendation in Kuaishou by utilizing cached results, monitored by a reinforcement learning-based framework. While most of the existing sequential methods leverage global information for personalized sequence modeling, sequence information is rarely explored for better global learning. We believe such unidirectional design limits the expressiveness of the learned embeddings, and a bidirectional information flow between different data modes is the key towards better heterogeneous interaction learning.
3. Preliminaries
Table 1. Symbols and notations.
Symbol Definition
$\mathcal{U},\mathcal{I}$ user set and item set
$\mathbf{x}_{j}^{(l)}$ $j$-th global feature of layer $l$
$\mathbf{s}_{t}^{(l)}$ sequence feature of layer $l$ at timestamp $t$
$m,n,k$ number of dense, sparse and sequence features
$T,d$ sequence length and embedding dimension
Table 1 summarizes the main symbols and notations used throughout the paper. We use bold uppercase letters for matrices (e.g., $\mathbf{X}$), bold lowercase letters for vectors (e.g., $\mathbf{x}$), and lowercase letters for scalars (e.g., $n$). The element at the $i$-th row and $j$-th column of a matrix $\mathbf{X}$ is denoted as $\mathbf{X}(i,j)$. The transpose of $\mathbf{X}$ is denoted by the superscript ${\scriptscriptstyle\mathsf{T}}$ (e.g., $\mathbf{X}^{\scriptscriptstyle\mathsf{T}}$). We use superscript $u$ to denote users, and subscripts $i$ and $t$ to denote item and timestamp, respectively (e.g., $y^{u}_{i_{t}}$). We use $\mathbf{x}_{i}^{(l)}$ to denote the $j$-th global feature of the $l$-th layer, and $\mathbf{s}_{t}^{(t)}$ to denote the sequence feature of the $l$-th layer at timestamp $t$. We consider the scenario with $m$ dense features, $n$ sparse features, and $k$ sequence features of length $T$. We use $d$ to denote the embedding dimension of the CTR model.
3.1. Click-Through Rate (CTR) Prediction
CTR prediction aims to estimate the probability of a user clicking on an item given heterogeneous information such as global context and behavior sequences. Formally, given a user set $\mathcal{U}$ and an item set $\mathcal{I}$, the interaction sequence of a user $u\in\mathcal{U}$ is defined as $S^{u}=[i^{u}_{1},i^{u}_{2},\dots,i^{u}_{T}]$, where $i^{u}_{t}\in\mathcal{I}$ is the item interacted at time step $t$. CTR prediction aims to estimate the probability of a user clicking on a new item $i^{u}_{T+1}$, denoted as $y^{u}_{i_{T+1}}$, given the historical interaction sequence $S^{u}$. Formally, we seek to learn a function $f:\mathcal{U}\times\mathcal{I}\times\mathcal{S}\rightarrow[0,1]$, where $\mathcal{S}$ is the set of all possible sequences, such that:
(1) $P(y^{u}_{i_{T+1}}=1|u,i_{T+1},S^{u};\theta)=f(u,i_{T+1},S^{u};\theta).$
To optimize the model, the cross-entropy loss, which measures the difference between predicted and groundtruth click probabilities, is commonly employed as the objective function. Since the formulation in Eq. (1) incorporates temporal dynamics in user behaviors, it potentially leads to better CTR predictions than non-sequential methods that ignore the sequential nature of user interactions.
3.2. Feature Interaction
Interaction learning, which aims to capture the complex relationships between different features, is the key towards the success of CTR prediction. We briefly introduce three prominent feature interaction modules: inner product, DCNv2 (Wang et al., 2021b) and DHEN (Zhang et al., 2022).
Inner Product-based Interaction. Given input feature vector $\mathbf{x}\in\mathbb{R}^{d}$, the inner product-based interaction, exemplified by Factorization Machines (FM) (Rendle, 2010), learns a latent vector $\mathbf{v}_{j}\in\mathbb{R}^{r}$ for each feature $j$, whose inner product $\langle\mathbf{v}_{j},\mathbf{v}_{k}\rangle$ describes the interaction strength between $\mathbf{x}(j)$ and $\mathbf{x}(k)$. Hence, the second-order feature interactions can be modeled as:
$f_{FM}(\mathbf{x})=\sum_{j=1}^{d}\sum_{k=j+1}^{d}\langle\mathbf{v}_{j},\mathbf{v}_{k}\rangle\mathbf{x}(j)\mathbf{x}(k)+\sum_{j=1}^{d}w_{j}\mathbf{x}(j)+w_{0}.$
Deep & Cross Network (DCNv2)  (Wang et al., 2021b). The DCNv2 model combines a cross network for explicit feature interactions with a deep neural network for implicit feature interactions. Given input feature vector $\mathbf{x}^{(0)}\in\mathbb{R}^{d}$, the $l$-th layer of the cross network models the second order interaction as follows:
$\mathbf{x}^{(l+1)}=\mathbf{x}^{(0)}\odot\left(\mathbf{w}^{(l)}\mathbf{x}^{(l)}+\mathbf{b}^{(l)}\right)+\mathbf{x}^{(l)}.$
By stacking multiple cross layers, DCNv2 is capable of modeling explicit high-order interactions. Besides, a deep network, which is essentially an MLP, is employed to model implicit interactions between features, and the final prediction is further obtained by combining the outputs of both networks.
Deep Hierarchical Ensemble Network (DHEN) (Zhang et al., 2022). DHEN leverages the strengths of heterogeneous interaction modules and learns a hierarchy of interactions of different orders. The key idea behind DHEN is to use a layered structure where each layer is an ensemble of multiple heterogeneous interaction modules.
Formally, given the concatenation of $m$ input features $\mathbf{X}^{(l)}\in\mathbb{R}^{d\times m}$, the output of the $l$-th DHEN layer is computed as:
$\mathbf{X}^{(l+1)}=\text{Norm}\left(\text{Ensemble}_{i=1}^{k}\text{Interaction}_{i}(\mathbf{X}^{(l)})+\text{ShortCut}(\mathbf{X}^{(l)})\right),$
where $\text{Norm}(\cdot)$ is a normalization function, $\text{Interaction}_{i}$ are different interaction modules that are further ensembled by $\text{Ensemble}_{i=1}^{k}(\cdot)$, e.g., summation and concatenation, and $\text{ShortCut}(\cdot)$ is an MLP serving as the residual connection for deep layer stacking.
By ensembling different modules (e.g., inner product, DCNv2, etc.), DHEN leverages their complementary strengths within a unified framework. Besides, by stacking multiple DHEN layers, high-order feature interactions can be captured.
3.3. Attention Mechanism
Attention mechanisms (Bahdanau et al., 2014; Vaswani et al., 2017) have become an integral part of sequence modeling, serving as the core component behind various model designs, such as Transformer (Vaswani et al., 2017) and BERT (Devlin et al., 2018). In this section, we briefly review attention mechanisms, including Multi-Head Attention and Pooling by Multihead Attention (PMA).
Multi-Head Attention (MHA) (Vaswani et al., 2017) allows sequence modeling without regard to distance between tokens (Vaswani et al., 2017). Given an input sequence $\mathbf{S}=[\mathbf{s}_{1},\dots,\mathbf{s}_{T}]$, where $\mathbf{s}_{t}\in\mathbb{R}^{d}$ is the embedding vector at timestamp $t$, the self-attention operation is defined as:
(2) $\vspace{-2pt}\text{Attn}(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{softmax}\left(\mathbf{QK}^{\scriptscriptstyle\mathsf{T}}/\sqrt{d_{k}}\right)\mathbf{V},\vspace{-2pt}$
where $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{n\times d_{k}}$ are query, key and value representations derived from the input sequence $\mathbf{S}$ via linear projections $\mathbf{Q}=\mathbf{S}^{\scriptscriptstyle\mathsf{T}}\mathbf{W}^{Q},\mathbf{K}=\mathbf{S}^{\scriptscriptstyle\mathsf{T}}\mathbf{W}^{K},\mathbf{V}=\mathbf{S}^{\scriptscriptstyle\mathsf{T}}\mathbf{W}^{V}$.
To enable joint attention to information from different embedding subspaces, MHA aggregates $h$ parallel attention as follows:
(3) $\vspace{-2pt}\text{MHA}(\mathbf{Q},\mathbf{K},\mathbf{V})=\left[\textbf{head}_{1}\|\ldots\|\textbf{head}_{h}\right]\mathbf{W}^{O},\vspace{-2pt}$
where heads are computed by self-attention $\textbf{head}_{i}=\text{Attn}(\mathbf{QW}_{i}^{Q},\allowbreak\mathbf{KW}_{i}^{K},\mathbf{VW}_{i}^{V})$ and aggregated via an output projector $\mathbf{W}^{O}\in\mathbb{R}^{hd_{k}\times d}$.
Pooling by Multi-Head Attention (PMA) (Lee et al., 2019). Rather than obtaining query, key and values from the same input sequence $\mathbf{S}$, PMA utilizes a learnable query $\mathbf{Q}_{\text{PMA}}\in\mathbb{R}^{k\times d_{k}}$ to summarize the sequence from $k$ different aspects, which can be defined as follows
(4) $\vspace{-2pt}\text{PMA}(\mathbf{Q}_{\text{PMA}},\mathbf{S})=\text{MHA}(\mathbf{Q}_{\text{PMA}},\mathbf{K},\mathbf{V}).\vspace{-2pt}$
Intuitively, each column in $\mathbf{Q}_{\text{PMA}}$ is a seed vector summarizing $\mathbf{S}$ as a $d$-dimensional vector, and the output of PMA is the concatenation of $k$ summarization depicting sequence from different aspects.
4. Methodology
Figure 1. An overview of the InterFormer model architecture. Each InterFormer block consists of three parts, including (1) Global Arch (orange) to learn behavior-aware global embeddings given sequence queries; (2) Sequence Arch (blue) to learn context-aware sequence embeddings given global queries; (3) Bridging Arch (green) to exchange information between global and sequence arch. Note that the dashed green line for CLS token appending only happens at the first layer.
In this section, we present our proposed InterFormer. We first introduce the preprocessing module in Section 4.1, followed by three major modules, including Global Arch, Sequence Arch and Bridging Arch. To learn behavior-aware global embeddings, the Global Arch is proposed to model feature interactions (Section 4.2). To learn context-aware sequence embeddings, the Sequence Arch is proposed to optimize sequence (Section 4.3). The Bridging Arch connects Global and Sequence Archs, enabling effective information summarization and exchange between different modes (Section 4.4). An overview of the InterFormer architecture is shown in Figure 1.
4.1. Feature Preprocessing
Global Feature Preproessing. We consider two types of global features, including dense features like user age and item price, and sparse features like user id and item category. To unify the heterogeneity in global features, the global preprocessing module transforms features into embeddings of the same dimensionality.
Specifically, raw dense features $x_{\text{dense}_{i}}^{(0)}$ are first concatenated to form a dense vector $\mathbf{x}_{\text{dense}}^{(0)}=\left[x_{\text{dense}_{1}}^{(0)},\dots,x_{\text{dense}_{m}}^{(0)}\right]^{\scriptscriptstyle\mathsf{T}}$, which is further transformed into a $d$-dimensional dense embedding vector via linear transformation, i.e., $\mathbf{x}_{\text{dense}}^{(1)}=\mathbf{W}_{\text{dense}}\mathbf{x}_{\text{dense}}^{(0)}$. Similarly, each raw sparse feature is first encoded as an one-hot vector $\mathbf{x}_{\text{sparse}_{i}}^{(0)}\in\mathbb{R}^{n_{v_{i}}}$, where $n_{v_{i}}$ is the vocabulary size of the $i$-th sparse feature, and further transformed into a $d$-dimensional embedding vector by $\mathbf{x}_{\text{sparse}}^{(1)}=\mathbf{W}_{\text{sparse}}\mathbf{x}_{\text{sparse}}^{(0)}$.
By concatenating the dense and sparse embedding vectors, the input global embedding matrix can be obtained as follows
(5) $\mathbf{X}^{(1)}=\left[\mathbf{x}_{\text{dense}}^{(1)}\|\mathbf{x}_{\text{sparse}_{1}}^{(1)}\|\dots\|\mathbf{x}_{\text{sparse}_{n}}^{(1)}\right].$
Sequence Feature Preprocessing. Similar to the global feature preprocessing, an embedding layer is employed such that each interacted item in the sequence is mapped to a $d$-dimensional vector $\mathbf{s}_{t}$, and the user behavior sequences is represented as the concatenation of item embeddings, i.e., $\mathbf{S}^{(0)}=\left[\mathbf{s}_{1}^{(0)}\|\dots\|\mathbf{s}_{T}^{(0)}\right]\in\mathbb{R}^{d\times T}$.
In real-world scenarios, we should expect multiple sequences originating from different user actions (e.g., click and conversion), different platforms, etc. Besides, due to the uncertainty of user behaviors, these sequences often contain noisy and irrelevant user-item interactions. To this end, a MaskNet (Wang et al., 2021a) is employed to unify multiple sequences and filter out the internal noises via self-masking.
Specifically, given $k$ sequences $\mathbf{S}_{1},\cdots,\mathbf{S}_{k}$, they are first concatenated along the embedding dimension, i.e., $\left[\mathbf{S}_{1}^{(0)^{\scriptscriptstyle\mathsf{T}}}\|\cdots\|\mathbf{S}_{k}^{(0)^{\scriptscriptstyle\mathsf{T}}}\right]^{\scriptscriptstyle\mathsf{T}}\in\mathbb{R}^{kd\times T}$, and further processed by the MaskNet operation as follows
(6) $\text{MaskNet}(\mathbf{S})=\text{MLP}_{\text{lce}}\left(\mathbf{S}\odot\text{MLP}_{\text{mask}}(\mathbf{S})\right),$
where $\text{MLP}_{\text{mask}}:\mathbb{R}^{kd\times T}\to\mathbb{R}^{kd\times T}$ generates the self-masking to select relevant information from the input sequence, and $\text{MLP}_{\text{lce}}:\mathbb{R}^{kd\times T}\to\mathbb{R}^{d\times T}$ linearly combines multiple sequences into one, matching the dimensionalities of global and sequence features.
4.2. Global Arch: Towards Behavior-Aware Interaction Learning
Global features, such as user profile and ad content, provide substantial information in understanding user preference over a specific item. Modeling the interaction among global features is the key towards the succuss of CTR prediction (Rendle, 2010; Wang et al., 2021b). While global features reflect static user interests (Lian et al., 2018; Zhang et al., 2022), the user behavior sequence provides complementary information depicting user interests from a dynamic view (Zhou et al., 2019). For instance, a user’s profile might indicate a general interest in electronics, while the recent browsing history on smartphones offers more specific and timely information about the current need. Therefore, it is crucial to learn behavior-aware global interactions to adapt to the evolving environment.
To learn behavior-aware global interaction, we model the interactions among global features, as well as sequence summarization, by an interaction module. Formally, given the global input $\mathbf{X}^{(l)}$ and sequence summarization $\mathbf{S}_{\text{sum}}^{(l)}$ at the $l$-th layer, the output of $l$-th Global Arch is defined as:
(7) $\mathbf{X}^{(l+1)}=\text{MLP}^{(l)}\left(\text{Interaction}^{(l)}\left(\left[\mathbf{X}^{(l)}\|\mathbf{S}_{\text{sum}}^{(l)}\right]\right)\right),\vspace{-5pt}$
where $\text{Interaction}^{(l)}(\cdot)$ is the interaction module. Note that we do not adhere to a specific interaction module, but rather, various backbone models, such as inner product, DCNv2 (Wang et al., 2021b) and DHEN (Zhang et al., 2022), can be adopted. Besides, an MLP is further employed to transform the output $\mathbf{X}^{(l+1)}$ to have consistent shape as the input $\mathbf{X}^{(l)}$ for selective information aggregation and easy layer-stacking.
By performing interaction learning on the concatenation of $\mathbf{X}^{(l)}$ and $\mathbf{S}^{(l)}_{\text{sum}}$, Eq. (7) incorporates three different interactions, including global-global, global-sequence, and sequence-sequence interactions. The global-global interactions capture explicit user interests by computing the relevance between the user profile and the target item content, while the global-sequence interactions capture implicit user interests by computing the relevance between the target item and the user’s current need behind the interaction sequence. Besides, by computing the interactions among sequence features, the most representative information can be promoted with high interaction scores, while inactive items with low scores can be filtered out. By stacking multiple layers, the Global Arch captures rich behavior-aware interactions at different orders.
4.3. Sequence Arch: Towards Context-Aware Sequence Modeling
In addition to the explicit static user preference in global features, the implicit dynamic user interests behind the behavior sequences provides complementary information (Zhou et al., 2019; Sun et al., 2019). However, given that user behavior sequence can be highly random and noisy, solely relying on sequence information is ineffective, and it is important to incorporate global context into sequence modeling. For example, a user may randomly browse items on online shopping platforms, but given the global information that the user is an electronic fan, electronic items, e.g., smartphone and laptops, in the browsing history are identified as key information require extra attention.
To learn context-aware sequence embeddings, a Sequence Arch is designed on top of two key ideas, including Personalized FFN (PFFN) and Multi-Head Attention (MHA). To enable interactions between global and sequence information, we first employ PFFN to transform sequence embeddings given global summarization as query. Given global summarization $\mathbf{X}_{\text{sum}}^{(l)}$ and sequence embedding $\mathbf{S}^{(l)}$ at the $l$-th layer, the PFFN operation is defined as:
(8) $\text{PFFN}\left(\mathbf{X}_{\text{sum}}^{(l)},\mathbf{S}^{(l)}\right)=f\left(\mathbf{X}_{\text{sum}}^{(l)}\right)\mathbf{S}^{(l)},\vspace{-5pt}$
where $f\left(\mathbf{X}_{\text{sum}}^{(l)}\right)$ is an MLP that aims to learn the linear projection on sequence based on global summarization.
Besides, to model the relationship among events in a sequence, MHA is applied to enable the model to attend to different parts of the sequence, capturing long-range dependencies and contextual information. To incorporate global context, before feeding into the first InterFormer layer, the global summarization $\mathbf{X}^{(1)}_{\text{sum}}$ is prepended before the sequence as the CLS token, i.e., $\mathbf{S}^{(1)}=\left[\mathbf{X}^{(1)}_{\text{sum}}\|\mathbf{S}^{(1)}\right]$, hence the following MHA on CLS token can aggregate sequence information given global information as query (Vaswani et al., 2017; Devlin et al., 2018). This is similar to the early fusion idea described in Transact  (Xia et al., 2023) but we primarily consider ‘append’ instead of ‘concat’ as the fusion method. Besides, rotary position embeddings (Su et al., 2024) are applied on tokens so that the positional information in sequences can be effectively leveraged. In general, the Sequence Arch can be written as follow:
(9) $\mathbf{S}^{(l+1)}=\text{MHA}^{(l)}\left(\text{PFFN}\left(\mathbf{X}_{\text{sum}}^{(l)},\mathbf{S}^{(l)}\right)\right).\vspace{-5pt}$
Since the output $\mathbf{S}^{(l+1)}$ is of the same shape as the input $\mathbf{S}^{(l)}$, aggressive sequence summarization can be avoided and layers can be easily stacked.
Through layer stacking, for one thing, the sequence embeddings is aware of global interactions at different orders via PFFN; for another, MHA at different layers focuses on different parts of the sequence, capturing multi-scale sequence information. Therefore, the model can learn rich context-aware encoding of the sequential data capturing both local and global patterns within the sequence.
4.4. Bridging Arch: Towards Effective Information Selection and Summarization
Though selective information aggregation is achieved in both the global and sequence arch, as the dimensionalities of input and output features are retained till the final layer, it is infeasible to directly exchange such information due to (1) ineffectiveness given the noisy information and (2) inefficiency given the high-dimensionality. To solve this dilemma, we propose a bridging arch to select and summarize information before exchanging between different data modes.
To start with, given the large scale of global embeddings, it is important to selectively summarize them to guide sequence learning. To this end, we highlight the most useful information through a personalized gated selection mechanism as follows:
(10) $\mathbf{X}_{\text{sum}}^{(l)}=\text{Gating}(\text{MLP}(\mathbf{X}^{(l)})),$
where $\text{where }\text{Gating}(\mathbf{X})=\mathbf{X}\odot\text{MLP}(\mathbf{X}).$
where MLP: $\mathbb{R}^{d\times n}\to\mathbb{R}^{d\times n_{\text{sum}}}$ with $n_{\text{sum}}\ll n$, and self-gating (Chai et al., 2020) provides sparse masking on the embeddings such that relevant information is retained while irrelevant noises are filtered out, providing high-quality context for efficient sequence learning.
For sequence information, three types of summarization are neatly designed, including CLS tokens, PMA tokens and recent interacted tokens. The CLS tokens $\mathbf{S}_{\text{CLS}}$ learned by MHA are selected as context-aware sequence summarization. However, the quality of CLS tokens largely depend on the learned global context. To compensate the heavy reliance on global context and enable more flexibility, the PMA tokens $\mathbf{S}_{\text{PMA}}$ (Lee et al., 2019), which are essentially sequence summarization based on learnable queries, are employed. Besides, the $K$ most recent interacted tokens $\mathbf{S}_{\text{recent}}$ are proven to be effective in capturing user’s recent interests (Borisyuk et al., 2024; Xia et al., 2023). The combination of the above information is further gated by a self-gating layer, serving as the behavior summarization for global interaction:
(11) $\mathbf{S}^{(l)}_{\text{sum}}=\text{Gating}\left(\left[\mathbf{S}^{(l)}_{\text{CLS}}\|\mathbf{S}^{(l)}_{\text{PMA}}\|\mathbf{S}^{(l)}_{\text{recent}}\right]\right).$
In general, the benefits of the bridging arch are two-fold. On the one hand, by separating information summarization from global and sequence arch, information can be retained in both arches, avoiding aggressive information aggregation. On the other hand, effective information exchange can be achieved as high-dimensional global/sequence features are selected and summarized into low-dimensional embeddings. Therefore, the Bridging Arch plays a pivotal role in enabling the model to capture complex interactions between global behaviors and sequential patterns, leading to more comprehensive representations of the input data.
5. Experiment
We evaluate the proposed InterFormer to answer the following research questions:
  • Q1: How effective is the proposed InterFormer (Section 5.2)?
  • Q2: How scalable is the proposed InterFormer (Section 5.3)?
  • Q3: To what extent does InterFormer address the insufficient inter-mode interaction and aggressive information aggregation (Section 5.4)?
  • Q4: To what extent does InterFormer benefit from different modules (Section 5.5)?
5.1. Experiment Setup
Datasets. We adopt three benchmark benchmark datasets, including AmazonElectronics (He and McAuley, 2016b), TaobaoAd (Tianchi, [n. d.]), KuaiVideo (Li et al., 2019), and a large-scale internal dataset for evaluation. Dataset statistics are summarized in Table 2 with more details in Appendix B.1.
Table 2. Dataset Summary.
Dataset #Samples #Feat. (Seq/Global) Seq Length
Amazon 3.0M 6 (2/4) 100
TaobaoAd 25.0M 22 (3/19) 50
KuaiVideo 13.7M 9 (4/5) 100
Baseline Methods. We compare the proposed InterFormer with 11 state-of-the-art models, including (1) non-sequential methods: FM (Rendle, 2010), xDeepFM (Lian et al., 2018), AutoInt+ (Song et al., 2019), DCNv2 (Wang et al., 2021b), FmFM (Sun et al., 2021), DOT product, DHEN (Zhang et al., 2022), Wukong (Zhang et al., 2024), and (2) sequential methods: DIN (Zhou et al., 2018), DIEN (Zhou et al., 2019), BST (Chen et al., 2019), DMIN (Xiao et al., 2020), DMR (Lyu et al., 2020). We adopt DHEN (Zhang et al., 2022) as the global arch for InterFormer in the experiments. Detailed model configurations are provided in Appendix B.2.
Metrics. We adopt three widely-used metrics to evaluate the models from different aspects, including:
  • AUC provides an aggregated measure of model capacity in correctly classifying positive and negative samples across all thresholds. Higher the better.
  • gAUC provides personalized AUC evaluation, where users are weighted by their click account. Higher the better.
  • LogLoss (cross-entropy loss) measures the distance between the model prediction $\hat{y}$ and the actual label $y$, and can be computed as $L(y,\hat{y})=-\left(y\log\left(\hat{y}\right)+(1-y)\log\left(1-\hat{y}\right)\right)$. Lower the better.
  • NE (normalized entropy) (He et al., 2014), is the LogLoss normalized by the entropy of the average empirical CTR of the training set. NE provides a data-insensitive evaluation on model performance as the normalization alleviates the effect of background CTR on model evaluation. Lower the better.
5.2. Effectiveness Results
Table 3. Experiments on benchmark datasets. Methods with high gAUC and AUC, and low LogLoss and #Params are preferred.
Dataset AmazonElectronics TaobaoAds KuaiVideo
Metric gAUC AUC LogLoss #Params gAUC AUC LogLoss #Params gAUC AUC LogLoss #Params
FM 0.8494 0.8485 0.5060 4.23M 0.5628 0.6231 0.1973 43.08M 0.6567 0.7389 0.4445 52.76M
xDeepFM 0.8763 0.8791 0.4394 5.46M 0.5675 0.6378 0.1960 53.79M 0.6621 0.7423 0.4382 43.56M
AutoInt+ 0.8786 0.8804 0.4441 5.87M 0.5701 0.6467 0.1941 42.05M 0.6619 0.7420 0.4369 43.95M
DCNv2 0.8783 0.8807 0.4447 5.23M 0.5704 0.6472 0.1933 43.71M 0.6627 0.7426 0.4378 42.48M
FmFM 0.8521 0.8537 0.4796 4.21M 0.5698 0.6330 0.1963 43.06M 0.6552 0.7389 0.4429 51.97M
DOT 0.8697 0.8703 0.4485 4.23M 0.5701 0.6482 0.1941 41.54M 0.6605 0.7435 0.4361 41.29M
DHEN 0.8759 0.8790 0.4398 4.99M 0.5708 0.6509 0.1929 43.89M 0.6589 0.7424 0.4365 42.06M
Wukong 0.8747 0.8765 0.4455 4.28M 0.5693 0.6478 0.1932 41.72M 0.6587 0.7423 0.4372 41.37M
DIN 0.8817 0.8848 0.4324 5.40M 0.5719 0.6507 0.1931 42.26M 0.6621 0.7437 0.4353 43.03M
[…]
BST 0.8823 0.8847 0.4305 5.30M 0.5698 0.6489 0.1935 42.05M 0.6617 0.7446 0.4352 42.83M
[…]
InterFormer 0.8843 0.8865 0.4253 7.18M 0.5728 0.6528 0.1926 44.73M 0.6637 0.7453 0.4340 43.61M
The experiment results are shown in Table 3. We first observe that sequential methods consistently outperforms non-sequential methods on all datasets. For non-sequential methods, sequence information is naively aggregated in early stages and further processed together with global information. While in sequential methods, neatly designed sequence processing modules, e.g., RNN (Zhou et al., 2019; Sun et al., 2019) and attention mechanism (Zhou et al., 2018; Chen et al., 2019; Xiao et al., 2017; Lyu et al., 2020), are employed, so that sequential information can be processed in aware of global context. The universal outperformance of sequential methods validates that different data modes should be processed differently. Besides, results show that aggressive sequence summarization in early stages in ignorance of global context will impair model performance.
Comparing InterFormer with other sequential methods, InterFormer achieves state-of-the-art performance. Specifically, InterFormer outperforms the best competitor by up to 0.14% in gAUC, 0.14% in AUC and 0.54% in LogLoss. These results demonstrate InterFormer’s effectiveness on diverse datasets and generalization across different recommendation tasks.
5.3. Scalability Results
To validate the scalability of InterFormer, we carry out experiments on a large-scale internal dataset containing 70B samples in total, hundreds of non-sequence features and 10 sequences of length 200 to 1,000. In general, a 3-layer InterFormer achieves a 0.15% NE gain compared to the internal SOTA model with similar FLOPs and 24% Queries Per Second (QPS) gain. Together with feature scaling, the improvement on NE can be further enlarged with a 10% of Model FLOPs Utilization (MFU) on 512 GPUs attained. InterFormer shows great generalizability on a wide range of models showing promising ROI.
On Sequence Feature Scaling. We evaluate how the performance of InterFormer changes when the sequence feature scale increases in internal dataset. In addition to six sequences of length 100, we include two additional long sequences of length 1000, and we observe a 0.14% improvement in NE. As shown in Figure 2(a), InterFormer exhibits better scalability compared to the strong internal baseline (Internal) that leverages cross-attention to capture sequence and non-sequence information, as NE curve of InterFormer continues to decrease when more training samples are involved, outperforming Internal by 0.06% in NE. Besides, we also tried to merge six sequences to generate one long sequence of length 600, and we observe promising efficiency improvements in QPS (+20%) and MFU (+17%) with a 0.02% NE tradeoff. From the modeling perspective, the results validate that InterFormer is able to enlarge the NE gain brought by the sequence feature scaling.
(a) Feature scaling
(b) Model scaling
Figure 2. Scalability of InterFormer on (a) feature scaling and (b) model scaling. InterFormer achieves better scaling curve than Internal, and promising layer-stacking performance on internal dataset.
On Model Scaling. Besides feature scaling, we evaluate how InterFormer performs when the model scale increases. As Figure 2(b) shows, scaling InterFormer from 1 to 4 layers achieves consistent NE gains, exhibiting good scaling properties. Specifically, compared to a single layer InterFormer, a two-layer InterFormer achieves a significant 0.13% NE gain, and 0.05% and 0.04% additional NE gains can be achieved when stacking 3 and 4 layers.
Model-System Co-Design. Model architecture design plays a crucial role in training efficiency with its implication on GPU FLOPs utilization and inter-GPU communication. We highlight two optimizations based on InterFormer architecture that in total boost training efficiency by more than 30%, including communication overhead reduction and computation efficiency.
For communication overhead reduction, the global arch DHEN with heavy parameters and relatively light computation tends to be FSDP ((Zhao et al., 2023)) communication bound in distributed training, while the sequence arch Transformer with much higher computation (FLOPs) to parameter ratio is normally computation bound. To alleviate such inefficiency, our parallel design of global and sequence arch allows the exposed communication from DHEN modules to overlap with sequence computation effectively, resulting in a 20% QPS improvement compared to the performing two modules sequentially.
For computation efficiency, we perform a series of optimizations, including 1) reallocating FLOPs from small, low-return modules to larger, high ROI modules and 2) combining smaller kernels to better utilize GPU resources. These optimizations improve MFU for interaction modules from 11% to 16%, and DHEN from 38% to 45%, with a 19% MFU improvement for the overall InterFormer layer.
5.4. Analysis of the Model
On the Interleaving Learning Style.
(a) DOT
(b) DCNv2
(c) DHEN
[…]
We first analyze how the interleaving learning style benefits heterogeneous interaction learning. To validate the universality of InterFormer, we consider three backbone global architectures, including dot product (DOT), DCNv2, and DHEN, and five different scenarios, including
  • sole where only global arch is adopted, while sequence information is naively aggregated in early stage.
  • sep where inter-mode interaction is disabled and sequence and global arch are learnt separately.
  • s2g where only sequence-to-global information flow is enabled, while the reverse direction is disabled.
  • g2s where only global-to-sequence information flow is enabled, while the reverse direction is disabled.
  • int where the bidirectional information flows are activated.
As shown in Figure 3, int consistently outperforms other scenarios regardless of the backbone global arch, with an up to 1.46% outperformance in AUC. This validates the universal benefits brought by our proposed interleaving learning style.
[…]
On the Selective Information Aggregation. We evaluate the effect of selective information aggregation by comparing with three aggressive aggregation variants, where sequence information is first compressed by (1) average pooling, (2) MLP, and (3) MHA, and further combined with global features by the interaction module DHEN. The experimental results are shown in Figure 4.
Figure 4. Study on information aggregation. We compare the performance with early summarization via average pooling, MLP, and MHA, and InterFormer’s selective aggregation.
As the results suggest, selective aggregation (InterFormer) consistently achieves the best performance compared with other aggressive variants. In general, we can regard MLP as a more selective version of average pooling as MLP gradually compresses the sequence information, and similarly, MHA process sequence less aggressive than MLP. Therefore, we may conclude that with more selective aggregation, the CTR prediction quality improves.
On Sequence Modeling. We also visualize the learned attention map in Figure 5 to understand sequence modeling.
Figure 5. Attention map on TaobaoAds. The first 4 tokens are CLS tokens, followed by the behavior sequence of length 50.
First, we observe that the attention map exhibits a cluster structure, i.e., surrounding tokens are likely to interact, which can be regarded as a weighted pooling within selected surrounding tokens. Such selective pooling process alleviates random noises in the behavior history to extract better temporal user interests.
Secondly, we find that MHAs at different layers focus on different tokens at different scales. For example, the attention map is more uniform in the first layer, indicating pooling in a broad scale for long-term interests. In contrast, the third layer exhibits stronger interactions within a smaller cluster, resulting in fine-grained aggregation for short-term interests. Furthermore, the second layer focuses on specific tokens, e.g., the recent tokens (tokens in right columns) and the 24-th token, for item-specific information.
5.5. Ablation Study
Figure 6. Ablation study on InterFormer, where ’-’ indicates ablating different information or modules.
To better understand how different model designs contribute to the final performance, we ablate different exchanged information, including PMA tokens and recent interacted tokens, and different modules, including the gating module, PFFN and MHA.
As the results shown in Figure 6, all the exchanged information and modules contribute to the model performance. Specifically, PMA tokens provide effective sequence summarization given global context and are of great importance, ablating which decreases AUC by up to 0.004. The gating modules help selective summarize information from different data modes contribute up to 0.003 AUC improvement. The sequence modeling modules (PFFN and MHA) also improves the performance to some extent.
6. Conclusion
In this paper, we propose a novel building block for heterogeneous interaction learning in CTR prediction named InterFormer. The key idea is an interleaving learning style between different data modes to achieve effective inter-mode interaction, as well as selective information aggregation. In general, A global arch and a sequence arch are adopted to achieve behavior-aware interaction learning and context-aware sequence modeling, respectively. A bridging arch is further proposed to for effective information selection and summarization. InterFormer achieves consistent outperformance on benchmark datasets, and 0.15% improvement on the NE of CTR prediction on internal large-scale dataset.
Appendix A Model Design
A.1. Algorithm
We present the overall algorithm of InterFormer in Algorithm 1.
Algorithm 1 InterFormer
0:  Global feat $\mathbf{X}^{(0)}$, sequence feat $\mathbf{S}^{(0)}$, number of layers $L$;
0:  CTR prediction $\hat{y}$;
1:  Preprocess $\mathbf{X}^{(0)},\mathbf{S}^{(0)}$ via Eqs. (5) and (6) to obtain $\mathbf{X}^{(1)},\mathbf{S}^{(1)}$;
2:  Compute global summarization $\mathbf{X}_{\text{sum}}^{(1)}$ via Eq. (10);
3:  Prepend $\mathbf{X}_{\text{sum}}^{(1)}$ before $\mathbf{S}^{(1)}$ as CLS tokens;
4:  for $l=1,2,\dots,L$ do
5:     Bridging Arch: compute global summarization $\mathbf{X}^{(l)}_{\text{sum}}$ and sequence summarization $\mathbf{S}^{(l)}_{\text{sum}}$ by Eqs. (10) and (11);
6:     Global Arch: compute global embeddings $\mathbf{X}^{(l+1)}$ given $\mathbf{X}^{(l)}$ and $\mathbf{S}^{(l)}_{\text{sum}}$ by Eq. (7);
7:     Sequence Arch: compute sequence embeddings $\mathbf{S}^{(l+1)}$ given $\mathbf{S}^{(l)}$ and $\mathbf{X}^{(l)}_{\text{sum}}$ by Eq. (9);
8:  end for
9:  Compute CTR prediction by $\hat{y}=\text{MLP}\left(\left[\mathbf{X}^{(L)}_{\text{sum}}\|\mathbf{S}^{(L)}_{\text{sum}}\right]\right)$;
10:  return  CTR prediction $\hat{y}$.
A.2. More Details on Module Blocks
We provide more details on module blocks involved in InterFormer architecture, including Linear Compressed Embedding (LCE), and personalized FeedForward Network (PFFN).
A.2.1. Linear Compressed Embedding
Given numerous global features, it is essential to compress embeddings to a more manageable size. In general, given $N$ $d$-dimensional features $\mathbf{X}\in\mathbb{R}^{d\times N}$, LCE is a linear transformation on the sample dimension $\mathbf{W}\in\mathbb{R}^{N\times M}$, such that the linear transformation $\mathbf{XW}\in\mathbb{R}^{d\times M}$ serves as the compressed embedding with $M$ features. Together with self-gating modules, LCE selectively compress and denoise numerous embeddings, benefiting the model in both efficiency and effectiveness.
A.2.2. Personalized FFN
PFFN learns interactions between global and sequence embeddings, whose key idea is to apply an FFN on the sequence embeddings with weight learned based on the summarized global embeddings. Given global summarization $\mathbf{X}_{\text{sum}}\in\mathbb{R}^{d\times n_{\text{sum}}}$ and sequence embedding $\mathbf{S}\in\mathbb{R}^{d\times T}$, PFFN first learns the transformation weight $\mathbf{W}_{\text{PFFN}}=\mathbf{X}_{\text{sum}}\mathbf{W}\in\mathbb{R}^{d\times d}$, which is further applied on the feature dimension of sequence embeddings, i.e., $\mathbf{W}_{\text{PFFN}}\mathbf{S}$. In contrast to the vanilla FFN in Transformer that may not add too much value to subsequent sequence modeling, PFFN integrate side information from other data modes. In fact, PFFN is essentially computing enhanced dot-products based interaction between sequence embeddings and summarized global embeddings that brings much more value for the following sequence modeling.
Appendix B Experiment Pipeline
We adopt the public BARS evaluation framework (Zhu et al., 2022) for benchmark experiments. In general, an Adam (Kingma, 2014) optimizer with a learning rate scheduler is adopted for optimization, where initial learning rate is tuned in {1e-1, 1e-2, 1e-3}. We use a batch size of 2048, and train up to 100 epochs with early stop. We adopt the Swish function (Ramachandran et al., 2017) for activation. We use NVIDIA A100 for benchmark experiments and NVIDIA H100 for internal experiments. More details on datasets and model configurations are provided as follows.
B.1. Datasets
We provide a brief description of datasets used in the paper.
  • AmazonElectronics (He and McAuley, 2016b) contains product reviews and metadata from Amazon with 192,403 users, 63,001 goods, 801 categories and 1,689,188 samples. Global features include user id, item id and item category. Sequence features include interacted items and corresponding categories with length of 100. We follow the public split with 2.60M samples for training and 0.38M samples for testing.
  • TaobaoAds (Tianchi, [n. d.]) contains 8 days of ad click-through data on Taobao (26 million records) randomly sampled from 1,140,000 users. Global features include item-related features, e.g., ad id, category, and price, and user-related features, e.g., user id, gender, and age. Sequence features include interacted items’ brands and categories, and user behaviors with length of 50. We follow the public split with 22.0M samples for training and 3.1M samples for testing.
  • KuaiVideo (Li et al., 2019) contains 10,000 users and their 3,239,534 interacted micro-videos. Global features include user id, video id, and visual embeddings of videos. Sequence features include different behaviors, e.g., click, like, and not click, with length of 100. We follow the public split with 10.9M samples for training and 2.7M samples for testing.
  • Internal contains 70B entries in total, with hundreds of global features describing users and items, and 10 sequence features of length 200 to 1,000.
B.2. Model Configuration
We provide the model configurations in the paper. We use the best searched model configurations on BARS whenever possible, and use the provided model default hyperparameters for the rest. In general, the attention MLP has sizes in {512,256,128,64} with number of heads ranged in {1,2,4,8}. An MLP with sizes in {1024,512,256,128} is adopted as the classifier head. Other model-specific configurations are provided as follows
  • xDeepFM (Lian et al., 2018): The compressed interaction network (CIN) is an MLP with size of 32.
  • DCNv2 (Wang et al., 2021b): The parallel structure is an MLP of size 512, and the stacked structure is an MLP with size of 500. We set the low-rank of the cross layer as 32.
  • DHEN (Zhang et al., 2022): We adopt DOT product and DCN as the ensembled modules. The number of layers ranged in {1,2,3}.
  • InterFormer: The number of cls tokens is set to be 4, the number of PMA tokens ranged is set to be 2, and the number of recent tokens ranged is set to be 2.